In a relay race, each contestant runs 25.0 m while carrying an egg balanced on a spoon, turns around, and comes back to the starting point. Edith runs the first 25.0 m in 20.0 s. On the return trip she is more confident and takes only 15.0 s. What is the magnitude of her average velocity for (a) the first 25.0 m? (b) The return trip? (c) What is her average velocity for the entire round trip? (d) What is her average speed for the round trip?

1
Expert's answer
2021-11-16T09:53:56-0500

(a) The average velocity is


"v=d/t=1.25\text{ m/s}."

(b) The average velocity on return trip:


"v=d/t=1.67\text{ m/s}."

(c) The average velocity here is total displacement over total time: (25-25)/(20+15)=0.

(d) The average speed is total distance over total time:


"s=\frac{25+25}{20+15}=1.43\text{ m/s}."